OpenStudy (jake124):

e + 7 = 15 A. 22 B. 8 my answer C. -8 D. -22 r + 18 = 26 A. 8 my answer B. 6 C. 44 D. 54 a - 6 = 19 A. 25 my answer B. 13 C. -13 D. -25 b - 3 = -4 A. 7 B. 1 my answer C. -1 D. -7

OpenStudy (virnios):

For the First one, your answer was correct

563blackghost (563blackghost):

The last one is incorrect but the rest are right.

OpenStudy (jake124):

ok is the last one -7 ?

Awolflover1 (awolflover1):

2nd one is correct because \[r+18−18=26−18\] 3rd is correct \[a−6+6=19+6\] last one is incorrect^ Add 3 to both sides. b−3+3=−4+3